carmelle:

There are 8 players on a tennis team. The team is planning to play in a doubles tournament. How many different groups of players of 2 players can the coach make, if the position does not matter? A. 28 B. 64 C. 20,160 D. 40,320

Arieonna:

i got 20,160 as my answer

Arieonna:

i will bring my explanation in a minute

Arieonna:

Using the arrangements formula, it is found that there are 20,160 different groups of players of 2 players that the coach can make. What is the arrangements formula? The number of possible arrangements of n elements is given by the factorial of n, that is: In this problem, all 8 players will play, hence considering the order, the number of ways is: 40,320/2 = 20,160.

Arieonna:

you can loom up the answer on brainly.com

carmelle:

lol u copied it straight from brainly, but that's okay, I figured out the answer already. thanks for the effort though! :)

Arieonna:

i am still trying to figure out my history problem i am having trouble on at the very top of the question center it is History i really need help on it

carmelle:

The answer is actually A. 28. Here's what I did, using the combination formula: \[\frac{ 8! }{ 2! \div 6! } = 28\]

carmelle:

@carmelle wrote:
The answer is actually A. 28. Here's what I did, using the combination formula: \[\frac{ 8! }{ 2! \div 6! } = 28\]
my bad, I meant: \[\frac{ 8! }{ 2! \times 6! } = 28\]